OpenStudy (anonymous):

The height of a triangle is 6 m more than its base. The area of the triangle is 56 m2. Find the length of the base. is it c? A. 7 m B. 8 m C. 14 m D. 15 m

OpenStudy (anonymous):

@animal108

OpenStudy (anonymous):

base^2+6*base-112=0 base^2+14*base-8*base-112=0 base*(base+14)-8*(base+14)=0 (base-8 )*( base+14)=0 base = 8 or base = -14 length can not be negative base = 8 hope this helped :)

OpenStudy (anonymous):

umm can you help me if i put a nother question up?

OpenStudy (anonymous):

yes i will help you

OpenStudy (anonymous):

yes tag me.